If c =25 billion,I= R35 billion and c =0.80
The equilibrium level of income is
1
Expert's answer
2021-07-14T12:03:07-0400

At equilibrium: AD = AS

Y = C + I + G

C = 25 + 0.80Y

I = 35

Y = C + I

Y = 25 + 0.80Y + 35

Y – 0.8Y = 25 + 35

0.2Y = 60

Y = "\frac{60}{0.2} = 300"


The equilibrium level of income (Y) = 300 billion.